OpenStudy (anonymous):

Find all real solutions of the equation. (Enter your answers as a comma-separated list. If there is no real solution, enter NO REAL SOLUTION.) (sqrt2x + 9)+ 3 = x

OpenStudy (anonymous):

rewrite question please

OpenStudy (anonymous):

Sqaure Root of 2x+9 +3=x

OpenStudy (anonymous):

\[\sqrt{2x+9}\]

OpenStudy (anonymous):

+3=x

OpenStudy (anonymous):

thanks NO SOLUTION because it defines itself so its an infinity loop

OpenStudy (anonymous):

the real solution is x = 1+sqrt(13)

OpenStudy (anonymous):

what about this

OpenStudy (anonymous):

think about the question at the very end you state the something is =x. Plug that expression into it self and you have a forever loop.

OpenStudy (anonymous):

x = -3

OpenStudy (anonymous):

no possible santistebanc because then the sqrt would have to be a negative

OpenStudy (anonymous):

\(\sqrt{2x+9}+3=x\).. Start by adding -3 to both sides and then square both sides \(\sqrt{2x+9}=x-3 \implies 2x+9=(x-3)^2 \implies 2x+9=x^2-6x+9\). Rearrange the terms after adding similar term you will get \(x^2-8x=0 \implies x(x-8)=0\). So, we have \(x=0\) or \(x=8\). If you check them by substituting in the original equation, you'll find that the only solution is \(x=8\).

OpenStudy (anonymous):

what did you mean? sqrt(2 x+9+3) = x or sqrt(2 x+9)+3 = x the first one is x = 1+sqrt(13) and the other is x = 8 both are real

OpenStudy (anonymous):

now that im so confused

OpenStudy (anonymous):

To be honest, you're the source of confusion because you didn't even write the question right. But, if the equation you want to solve is what I wrote in the first line of my comment, then you just have to follow the solution I gave.

OpenStudy (anonymous):

yeah agree with AnwarA